# MiniPLM: Knowledge Distillation for Pre-Training Language Models

### v2

@@ Abstract
Knowledge distillation (KD) is widely used to train small, high-performing student language models (LMs) using large teacher LMs. While effective in fine-tuning, KD during pre-training faces challenges in efficiency, flexibility, and effectiveness. Existing methods either incur high computational costs due to online teacher inference, require tokenization matching between teacher and student LMs, or risk losing the difficulty and diversity of the teacher-generated training data. To address these issues, we propose MiniPLM, a KD framework for pre-training LMs by refining the training data distribution with the teacher’s knowledge. For efficiency, MiniPLM performs offline teacher LM inference, allowing KD for multiple student LMs without adding training-time costs. For flexibility, MiniPLM operates solely on the training corpus, enabling KD across model families. For effectiveness, MiniPLM leverages the differences between large and small LMs to enhance the difficulty and diversity of the training data, helping student LMs acquire versatile and sophisticated knowledge. Extensive experiments demonstrate that MiniPLM boosts the student LMs’ performance on 9 widely used downstream tasks, improves the language modeling capabilities, and reduces pre-training computation. The benefit of MiniPLM extends to large pre-training scales, evidenced by the extrapolation of the scaling curves. Further analysis reveals that MiniPLM supports KD across model families and enhances the utilization of pre-training data. Our model, code, and data are available at https://github.com/thu-coai/MiniPLM.
@@ 1 Introduction
Recent advances in language models (LMs; 35, 8, 60, 78) have largely been driven by scaling up model sizes, but this comes with high inference costs for deployment. At the same time, training small, deployment-friendly LMs faces training computation challenges, as small models are typically far from compute-optimal configurations according to Scaling Laws [37]. This has spurred a growing interest in exploring the limits of small LMs under the constraint of pre-training computation [51].
@@
Knowledge Distillation (KD; 36), where a small student model learns from a large teacher model, is a promising approach for training high-performing small LMs. While KD is effective for fine-tuning LMs on specific tasks [88], its role in improving pre-training, the critical stage for small LMs to acquire versatile foundation knowledge, remains under-explored. It is non-trivial to apply KD during pre-training with the methods in fine-tuning stages, which can be categorized as online KD and offline KD. In online KD, the teacher LM has to perform inference during the student LM pre-training to provide token-level probability supervision [^2] Pre-computing this supervision is impractical, as it requires 30PB of storage for 50B tokens with a 150K vocabulary. Therefore, teacher LM’s probabilities are typically computed online during the student LM training.  [65, 32], introducing additional training-time computation overhead. As shown in Figure 2, while online KD enhances performance within the same training steps, its benefits diminish if the extra computation was instead used to extend pre-training without KD. In addition, most online KD methods require the teacher and student LMs to share tokenization, limiting its flexibility across different model families. Offline KD, on the other hand, avoids extra training-time computation and allows for KD across model families, as student LMs are trained on the data offline generated by the teacher LM [41, 33]. However, ensuring sufficient difficulty and diversity in the teacher-generated data for pre-training is challenging without extensive human expertise. This often causes the student LM to overfit easy and common language patterns, hindering its generalization to versatile downstream tasks [70], as demonstrated in Figure 2.
@@
To address these challenges, we propose MiniPLM, an efficient, flexible, and effective KD framework for pre-training LMs, as illustrated in Figure 3. Intuitively, MiniPLM distills the teacher LM’s knowledge into the pre-training distribution through Difference Sampling, which samples training instances based on the “difference” between large and small LMs. Student LMs are then pre-trained from scratch on the refined distribution. To ensure efficiency, as shown in Figure 3(a), Difference Sampling performs offline teacher LM inference, allowing MiniPLM to distill knowledge into multiple student LMs without incurring additional training-time costs. For flexibility, MiniPLM operates solely on the training corpus, enabling KD across model families and ensuring seamless integration with highly optimized training pipelines [15]. In terms of effectiveness, Difference Sampling samples training instances that the teacher LM prefers but that a small reference LM assigns low probabilities to, promoting data difficulty and diversity. As depicted in Figure 3(b), this design down-samples easy and common patterns, up-samples hard and diverse instances, and filters out noisy or harmful data points from the pre-training corpus, which encourages student LMs to acquire versatile and sophisticated knowledge, ultimately improving downstream generalization.
@@
We apply MiniPLM to pre-train 200M, 500M, and 1.2B student LMs from scratch, using a 1.8B teacher LM. We show that MiniPLM surpasses various baselines in improving student LMs’ zero-shot performance on 9 widely used downstream tasks, enhancing language modeling capabilities, and reducing pre-training computation. By extrapolating the test loss with the Scaling Law [37], we observe that MiniPLM’s benefit remains consistent for LMs trained on ~10T tokens. MiniPLM also facilitates KD across model families, improving Llama3.1 [19] and Mamba [27] with a teacher LM from the Qwen family [3]. Further analysis shows that MiniPLM enhances pre-training data utilization, reducing the data demand by 2.4 times, which mitigates the quick exhaustion of web-crawled corpora [80].
@@ 2 MiniPLM: KD for Pre-training LMs
We consider pre-training an LM with an output distribution $q_{\bm{\theta}}$ on a large-scale corpus $\mathcal{D}$ consisting of $N$ text sequences, where ${\bm{\theta}}$ represents the model parameters. KD aids pre-training by incorporating the knowledge of a teacher LM with output distribution $p$, and training ${\bm{\theta}}$ to minimize the discrepancy between $p$ and $q_{\bm{\theta}}$ [65]. MiniPLM formulates KD as a reward maximization problem [34], which trains the student LM to generate diverse texts receiving high preference from the teacher LM (Section 2.1). To efficiently and effectively optimize the reward, as illustrated by Figure 3, MiniPLM employs Difference Sampling to refine the pre-training corpus (Section 2.2). This process improves the pre-training distribution in an offline manner with the teacher LM’s knowledge and preserves the data diversity and difficulty with the help of a small reference model. The student LM is then pre-trained from scratch on the refined corpus (Section 2.3).
@@ 2.1 KD as Reward Maximization
Recent works [32, 1, 42] have shown the effectiveness of minimizing the reverse Kullback-Leibler divergence (KLD) between $p$ and $q_{\bm{\theta}}$ for KD of LMs in the fine-tuning stage, which avoids $q_{\bm{\theta}}$ from over-estimating the low-probability regions of $p$. We reformulate this objective as a reward maximizing problem [34], which is suitable for designing efficient offline optimization methods [44]:
@@
where the reward $r$ is defined as $r(p,q_{\bm{\theta}},{\bm{x}})=\log\frac{p({\bm{x}})}{q_{\bm{\theta}}({\bm{x}})}$. Intuitively, Eq. (1) trains $q_{\bm{\theta}}$ to generate text ${\bm{x}}$ that receives high reward values, indicating a preference from the teacher LM (i.e., high $\log p({\bm{x}})$ values), while also ensuring high diversity (i.e., low $\mathbb{E}_{{\bm{x}}\sim q_{{\bm{\theta}}}}q_{\bm{\theta}}({\bm{x}})$ values). To optimize Eq. (1), a simple yet effective approach is Best-of-N [71, 4, 79], where a set $\mathcal{D}_{q_{\bm{\theta}}}$ containing $M$ candidates is first sampled from $q_{\bm{\theta}}$ and $K$ instances with the highest rewards are selected from these candidates to form a new dataset $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$:
@@
where $\mathcal{D}_{q_{\bm{\theta}}}=\{\left.x_{m}\right|x_{m}\sim q_{\bm{\theta}},1\leq m\leq M\}$. The student LM is then trained on $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$ to learn to generate texts with large $r(p,q_{\bm{\theta}},{\bm{x}})$ values. Although Best-of-N achieves performing the teacher LM’s inference prior to the student LM training similar to offline KD [41, 33], it still lacks the efficiency advantage of these methods because obtaining $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$ requires (1) sampling data from $q_{\bm{\theta}}$ and (2) computing the reward values with $q_{\bm{\theta}}$, making $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$ non-transferable for pre-training other student LMs. It is also hard to ensure the diversity and difficulty of the $M$ candidates sampled from $q_{\bm{\theta}}$ without careful prompt engineering with human expertise. In the following, we show that these issues can be effectively and efficiently addressed by Difference Sampling, which is the basis of the MiniPLM training algorithm.
@@ 2.2 Difference Sampling
The proof of Proposition 3 is provided in Appendix A. In our context, $p_{1}$ represents the data distribution of $\mathcal{D}$, $p_{2}=q_{\bm{\theta}}$, which is the data sampling distribution of $\mathcal{D}_{q_{\bm{\theta}}}$, and $r({\bm{x}})=r(p,q_{\bm{\theta}},{\bm{x}})$. Intuitively, Proposition 3 reveals the fact that when $|\mathcal{D}|=N$ and $|\mathcal{D}_{q_{\bm{\theta}}}|=M$ are sufficiently large, the top-$K$ instances selected from both sets tend to overlap. This is well-suited to our scenario, as the pre-training corpus $\mathcal{D}$ is typically large-scale, and it is advantageous to sample as many candidates from $\mathcal{D}_{q_{\bm{\theta}}}$ as possible to find high-reward instances.
@@
To avoid computing reward values with $\bm{q_{\bm{\theta}}}$, we replace it with $p_{\text{ref}}$, the output distribution of a tiny reference LM, typically smaller than the student LM, for reward computation. The reference LM is pre-trained on a small subset $\mathcal{D}_{\text{ref}}$, uniformly sampled from $\mathcal{D}$, allowing $p_{\text{ref}}$ to approximate $q_{\bm{\theta}}$ with minimal computation. This is a reasonable approximation because $q_{\bm{\theta}}$ evaluates the difficulty of instances, and the relative data difficulties generally remain consistent across different models [21]. As a result, the reward function in Eq. (2) is replaced with $r(p,p_{\text{ref}},{\bm{x}})=\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}$.
@@ 2.5 Discussion
As shown in Figure 3(a), MiniPLM relies only on $p({\bm{x}})$ and $p_{\text{ref}}({\bm{x}})$, the teacher and reference LMs’ probability over the entire sequence, which can be computed and stored offline because each instance in $\mathcal{D}$ is associated with a single floating-point number, amounting to only 200MB storage for 50B tokens with a 1,024 sequence length. Once $\mathcal{D}^{\prime}$ is difference-sampled based on $p({\bm{x}})$ and $p_{\text{ref}}({\bm{x}})$, multiple student LMs can be efficiently pre-trained under the teacher LM’s guidance without extra computational cost. In addition, this process modifies only the pre-training data, imposing no restrictions on the architecture or tokenization of the student LM, making MiniPLM highly flexible for integration into optimized pre-training frameworks [15] and suitable for KD across model families. In contrast, online KD [58, 32] relies on per-token distributions, which are infeasible to store offline, as it takes $50B\times 150K\times 4\text{byte}=30\text{PB}$ storage for an LM with 150K vocabulary trained on 50B tokens, making online teacher LM inference necessary. Aligning per-token distributions also demands matching tokenizers between the teacher and student [7], which complicates the re-implementation and optimization of pre-training workflows.
@@
In essence, as illustrated in Figure 3(b), MiniPLM distills the teacher LM’s knowledge into the student LM’s pre-training distribution via Difference Sampling, producing three effects: (1) Down-sampling easy and common patterns that both the teacher and reference LM fit well, where $p({\bm{x}})\gtrsim p_{\text{ref}}({\bm{x}})$ and $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}\gtrsim 0$. (2) Up-sampling hard and diverse knowledge, which the larger teacher LM has mastered but the smaller reference LM struggles with, and thus $p({\bm{x}})\gg p_{\text{ref}}({\bm{x}})$ and $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}\gg 0$. (3) Discarding noisy and harmful instances that the teacher LM assigns lower probabilities to than the reference LM, where $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}<0$. We provide examples of these effects in Appendix C.5. Training on the distribution with these effects encourages the student model to focus more on the sophisticated world knowledge learned by the teacher without being distracted by the noise. Note that without the comparison between $p$ and $p_{\text{ref}}$, the effect (1) disappears, leading to a pre-training corpus dominated by common patterns. This explains the effectiveness of MiniPLM against prior offline KD methods [41, 63], where maintaining the data difficulty and diversity, critical to pre-training [70], is challenging without extensive human efforts on prompt engineering [33].

### v3

@@ Abstract
Knowledge distillation (KD) is widely used to train small, high-performing student language models (LMs) using large teacher LMs. While effective in fine-tuning, KD during pre-training faces efficiency, flexibility, and effectiveness issues. Existing methods either incur high computational costs due to online teacher inference, require tokenization matching between teacher and student LMs, or risk losing the difficulty and diversity of the teacher-generated training data. In this work, we propose MiniPLM, a KD framework for pre-training LMs by refining the training data distribution with the teacher LM’s knowledge. For efficiency, MiniPLM performs offline teacher inference, allowing KD for multiple student LMs without adding training costs. For flexibility, MiniPLM operates solely on the training corpus, enabling KD across model families. For effectiveness, MiniPLM leverages the differences between large and small LMs to enhance the training data difficulty and diversity, helping student LMs acquire versatile and sophisticated knowledge. Extensive experiments demonstrate that MiniPLM boosts the student LMs’ performance on 9 common downstream tasks, improves language modeling capabilities, and reduces pre-training computation. The benefit of MiniPLM extends to larger training scales, evidenced by the scaling curve extrapolation. Further analysis reveals that MiniPLM supports KD across model families and enhances the pre-training data utilization. Our code, data, and models can be found at https://github.com/thu-coai/MiniPLM.
@@ 1 Introduction
Recent advances in language models (LMs; Han et al., 2021; Bommasani et al., 2021; OpenAI, 2023; Touvron et al., 2023a) have largely been driven by scaling up model sizes, but this comes with high inference costs for deployment. At the same time, training small, deployment-friendly LMs faces training computation challenges, as small models are typically far from compute-optimal configurations according to Scaling Laws (Hoffmann et al., 2022). This has spurred a growing interest in exploring the limits of small LMs under the constraint of pre-training computation (Lu et al., 2024).
@@
Knowledge Distillation (KD; Hinton et al., 2015), where a small student LM learns from a large teacher LM, is a promising approach for training high-performing small LMs. While KD is effective for fine-tuning (Xu et al., 2024), its role in improving pre-training, the critical stage for LMs to acquire foundation knowledge, remains under-explored. Applying KD during pre-training with the methods in fine-tuning stages is non-trivial, which can be categorized as online KD and offline KD. In online KD, the teacher LM has to perform inference during pre-training to provide token-level probability supervision [^2] Pre-computing this supervision is impractical, as it requires 30PB of storage for 50B tokens with a 150K vocabulary. Therefore, teacher LM’s probabilities are typically computed online during the student LM training.  (Sanh et al., 2019; Gu et al., 2024b), introducing additional training-time overhead. As shown in Figure 2, while online KD enhances performance within the same training steps, its benefits diminish if the extra computation was instead used to extend pre-training without KD. In addition, most online KD methods require the teacher and student LMs to share tokenization, limiting their flexibility for knowledge transfer across model families. Offline KD, on the other hand, avoids extra training-time computation and allows for KD across model families, as student LMs are trained on the data offline generated by the teacher (Kim & Rush, 2016; Gunasekar et al., 2023). However, ensuring sufficient difficulty and diversity in the generated data is challenging without extensive human expertise. Student LMs tend to overfit easy and common language patterns, hindering their downstream generalization (Shumailov et al., 2024), as shown in Figure 2.
@@
To address these challenges, we propose MiniPLM, an efficient, flexible, and effective KD framework for pre-training LMs, as illustrated in Figure 3. MiniPLM typically works in scenarios where the teacher LM’s probabilities are available to the users. Intuitively, it distills the teacher LM’s knowledge into the pre-training distribution through Difference Sampling, which samples training instances based on the “difference” between large and small LMs. Student LMs are then pre-trained from scratch on the refined distribution. To ensure efficiency, as shown in Figure 3(a), Difference Sampling performs offline teacher LM inference, allowing MiniPLM to distill knowledge into multiple student LMs without incurring additional training-time costs. For flexibility, MiniPLM operates solely on the training corpus, enabling KD across model families and ensuring seamless integration with highly optimized training pipelines (Chowdhery et al., 2022). In terms of effectiveness, Difference Sampling samples training instances that the teacher LM prefers but that a small reference LM assigns low probabilities to, promoting data difficulty and diversity. As depicted in Figure 3(b), this design down-samples easy and common patterns, up-samples hard and diverse instances, and filters out noisy or harmful data points from the pre-training corpus, which encourages student LMs to acquire versatile and sophisticated knowledge, ultimately improving downstream generalization.
@@
We apply MiniPLM to pre-train 200M, 500M, and 1.2B student LMs from scratch, using a 1.8B teacher LM. We show that MiniPLM surpasses various baselines in improving student LMs’ zero-shot performance on 9 widely used downstream tasks, enhancing language modeling capabilities, and reducing pre-training computation. By extrapolating the test loss with the Scaling Law (Hoffmann et al., 2022), we observe that MiniPLM’s benefit remains consistent for LMs trained on ~10T tokens. MiniPLM also facilitates KD across model families, improving Llama3.1 (Dubey et al., 2024) and Mamba (Gu & Dao, 2023) with a teacher LM from the Qwen family (Bai et al., 2023). Further analysis shows that MiniPLM enhances pre-training data utilization, reducing the data demand by 2.4 times, which mitigates the quick exhaustion of web-crawled corpora (Villalobos et al., 2022).
@@ 2 MiniPLM: KD for Pre-training LMs
We consider pre-training an LM with an output distribution $q_{\bm{\theta}}$ on a large-scale corpus $\mathcal{D}$ consisting of $N$ text sequences, where ${\bm{\theta}}$ represents the model parameters. KD aids pre-training by incorporating the knowledge of a teacher LM with output distribution $p$, and training ${\bm{\theta}}$ to minimize the discrepancy between $p$ and $q_{\bm{\theta}}$ (Sanh et al., 2019). MiniPLM formulates KD as a reward maximization problem (Haarnoja et al., 2017), which trains the student LM to generate diverse texts receiving high preference from the teacher LM (Section 2.1). To efficiently and effectively optimize the reward, as illustrated by Figure 3, MiniPLM employs Difference Sampling to refine the pre-training corpus (Section 2.2). This process improves the pre-training distribution in an offline manner with the teacher LM’s knowledge and preserves the data diversity and difficulty with the help of a small reference model. The student LM is then pre-trained from scratch on the refined corpus (Section 2.3).
@@ 2.1 KD as Reward Maximization
Recent works (Gu et al., 2024b; Agarwal et al., 2024; Ko et al., 2024) have shown the effectiveness of minimizing the reverse Kullback-Leibler divergence (KLD) between $p$ and $q_{\bm{\theta}}$ for KD of LMs in the fine-tuning stage, which avoids $q_{\bm{\theta}}$ from over-estimating the low-probability regions of $p$. We reformulate this objective as a reward maximizing problem (Haarnoja et al., 2017), which is suitable for designing efficient offline optimization methods (Levine et al., 2020):
@@
where the reward $r$ is defined as $r(p,q_{\bm{\theta}},{\bm{x}})=\log\frac{p({\bm{x}})}{q_{\bm{\theta}}({\bm{x}})}$. Intuitively, Eq. (1) trains $q_{\bm{\theta}}$ to generate text ${\bm{x}}$ that receives high reward values, indicating a preference from the teacher LM (i.e., high $\log p({\bm{x}})$ values), while also ensuring high diversity (i.e., low $\mathbb{E}_{{\bm{x}}\sim q_{{\bm{\theta}}}}q_{\bm{\theta}}({\bm{x}})$ values). To optimize Eq. (1), a simple yet effective approach is Best-of-N (Stiennon et al., 2020; Bai et al., 2022; Touvron et al., 2023b), where a set $\mathcal{D}_{q_{\bm{\theta}}}$ containing $M$ candidates is first sampled from $q_{\bm{\theta}}$ and $K$ instances with the highest rewards are selected from these candidates to form a new dataset $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$:
@@
where $\mathcal{D}_{q_{\bm{\theta}}}=\{\left.x_{m}\right|x_{m}\sim q_{\bm{\theta}},1\leq m\leq M\}$. The student LM is then trained on $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$ to learn to generate texts with large $r(p,q_{\bm{\theta}},{\bm{x}})$ values. Although Best-of-N achieves performing the teacher LM’s inference prior to the student LM training similar to offline KD (Kim & Rush, 2016; Gunasekar et al., 2023), it still lacks the efficiency advantage of these methods because obtaining $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$ requires (1) sampling data from $q_{\bm{\theta}}$ and (2) computing the reward values with $q_{\bm{\theta}}$, making $\mathcal{D}^{\prime}_{q_{\bm{\theta}}}$ non-transferable for pre-training other student LMs. It is also hard to ensure the diversity and difficulty of the $M$ candidates sampled from $q_{\bm{\theta}}$ without careful prompt engineering with human expertise. In the following, we show that these issues can be effectively and efficiently addressed by Difference Sampling, which is the basis of the MiniPLM training algorithm.
@@ 2.2 Difference Sampling
The proof of Proposition 3 is provided in Appendix A. Here, $p_{1}$ represents the data distribution of $\mathcal{D}$, $p_{2}=q_{\bm{\theta}}$, which is the sampling distribution of $\mathcal{D}_{q_{\bm{\theta}}}$, and $r({\bm{x}})=r(p,q_{\bm{\theta}},{\bm{x}})$. Intuitively, Proposition 3 reveals that when $|\mathcal{D}|=N$ and $|\mathcal{D}_{q_{\bm{\theta}}}|=M$ are sufficiently large, the top-$K$ instances selected from both sets tend to overlap. This suits our scenario well, as $\mathcal{D}$ is typically large-scale, and it is advantageous to sample as many candidates from $\mathcal{D}_{q_{\bm{\theta}}}$ as possible to find high-reward instances.
@@
To avoid computing reward values with $\bm{q_{\bm{\theta}}}$, we replace it with $p_{\text{ref}}$, the output distribution of a tiny reference LM, typically smaller than the student LM, for reward computation. The reference LM is pre-trained on a small subset $\mathcal{D}_{\text{ref}}$, uniformly sampled from $\mathcal{D}$, allowing $p_{\text{ref}}$ to approximate $q_{\bm{\theta}}$ with minimal computation. This is a reasonable approximation because $q_{\bm{\theta}}$ evaluates the difficulty of instances, and the relative data difficulties generally remain consistent across different models (Ethayarajh et al., 2022). As a result, the reward function in Eq. (2) is replaced with $r(p,p_{\text{ref}},{\bm{x}})=\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}$. In Appendix D.4, we empirically show that $r(p,p_{\text{ref}},{\bm{x}})$ and $r(p,q_{\text{ref}},\bm{x})$ have high correlations.
@@ 2.5 Discussion
As shown in Figure 3(a), MiniPLM relies only on $p({\bm{x}})$ and $p_{\text{ref}}({\bm{x}})$, the teacher and reference LMs’ probability over the entire sequence, which can be computed and stored offline because each instance in $\mathcal{D}$ is associated with a single floating-point number, amounting to only 200MB storage for 50B tokens with a 1,024 sequence length. Once $\mathcal{D}^{\prime}$ is difference-sampled based on $p({\bm{x}})$ and $p_{\text{ref}}({\bm{x}})$, multiple student LMs can be efficiently pre-trained under the teacher LM’s guidance without extra computational cost. In addition, this process modifies only the pre-training data, imposing no restrictions on the architecture or tokenization of the student LM, making MiniPLM highly flexible for integration into optimized pre-training frameworks (Chowdhery et al., 2022) and suitable for KD across model families. In contrast, online KD (Muralidharan et al., 2024; Gu et al., 2024b) relies on per-token distributions, which are infeasible to store offline, as it takes $50B\times 150K\times 4\text{byte}=30\text{PB}$ storage for an LM with 150K vocabulary trained on 50B tokens, making online teacher LM inference necessary. Aligning per-token distributions also demands matching tokenizers between the teacher and student (Boizard et al., 2024), which complicates the re-implementation and optimization of pre-training workflows.
@@
In essence, as illustrated in Figure 3(b), MiniPLM distills the teacher LM’s knowledge into the student LM’s pre-training distribution via Difference Sampling, producing three effects: (1) Down-sampling easy and common patterns that both the teacher and reference LM fit well, where $p({\bm{x}})\gtrsim p_{\text{ref}}({\bm{x}})$ and $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}\gtrsim 0$. (2) Up-sampling hard and diverse knowledge, which the larger teacher LM has mastered but the smaller reference LM struggles with, and thus $p({\bm{x}})\gg p_{\text{ref}}({\bm{x}})$ and $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}\gg 0$. (3) Discarding noisy and harmful instances that the teacher LM assigns lower probabilities to than the reference LM, where $\log\frac{p({\bm{x}})}{p_{\text{ref}}({\bm{x}})}<0$. We provide examples of these effects in Appendix F. Training on the distribution with these effects encourages the student model to focus more on the sophisticated world knowledge learned by the teacher without being distracted by the noise. Note that without the comparison between $p$ and $p_{\text{ref}}$, the effect (1) disappears, leading to a pre-training corpus dominated by common patterns. This explains the effectiveness of MiniPLM against prior offline KD methods (Kim & Rush, 2016; Peng et al., 2023), where maintaining the data difficulty and diversity, critical to pre-training (Shumailov et al., 2024), is challenging without extensive human efforts on prompt engineering (Gunasekar et al., 2023).